# MMBind: Unleashing the Potential of Distributed and Heterogeneous Data for Multimodal Learning in IoT

### v1

@@ 5.1. Pairing Incomplete Data with Shared Modality
Instead of directly comparing the high-dimensional raw sensor data, we propose to use a feature encoder to compress the shared modality data into a latent space and measure the similarity of the resulting embeddings. Recent works (Connor et al., 2021; Kirchoff et al., 2024; Xu et al., [n. d.]) show that properly trained autoencoders exhibit structured latent spaces, where semantically similar samples are grouped close together while dissimilar samples are further apart.
@@
Similarly, when labels serve as the shared modality, we can directly utilize pre-trained large language models (LLMs) to measure the semantic similarity of natural language labels, eliminating the need to train unimodal encoders. In this context, we employ an encoder-only LLM (paraphrase-MiniLM-L6-v2 trained by Sentence-BERT (Reimers, 2019)), which represents input text in a dense vector space for similarity measurement. In this paper, we primarily focus on cases where labels sufficiently describe the similarity of the data. However, class labels can sometimes fail to fully capture the similarity of an event, leading to incongruent data associations. For example, sensor data labeled identically but collected from different locations (e.g., worn inwards versus outwards) may exhibit significant differences. In such cases, we can enhance the input for language models by incorporating additional context or metadata in structured formats, such as the subjects’ gender, environment, and devices used, alongside the class labels. The LLM then identifies the most semantically similar samples and calculates their corresponding similarities based on the language tokens of the labels.
@@ 5.1.2. Generating Pseudo-Paired Data
As mentioned earlier, when labels serve as the shared modality, we utilize existing pre-trained LLMs to tokenize the language labels and derive their corresponding embeddings $\mathbf{h}^{j}_{s}$ and $\mathbf{h}^{k}_{s}$. Next, we compute the pairwise cosine similarity $a_{jk}$ between features of the shared modality $\mathbf{h}^{j}_{s}$ from $D_{A}$ and $\mathbf{h}^{k}_{s}$ from $D_{B}$, resulting in a similarity matrix $\mathbf{A}\in\mathbb{R}^{|D_{A}|\times|D_{B}|}$.
@@
Here, $\arg\max$ denotes the index where $a_{jk}$ achieves its maximum value within the array $\mathbf{a}_{j}$. Therefore, the corresponding paired sample from $D_{B}$ will be $\mathbf{X}^{p_{j}}_{B}=(\mathbf{m}^{p_{j}}_{2},\mathbf{m}^{p_{j}}_{s})$. Moreover, each pseudo-paired multimodal data sample will also be associated with the corresponding similarity measurement $a_{jp_{j}}$. This process occurs for every $j_{th}$ data sample in $D_{A}$ and again for every $k_{th}$ data sample in $D_{B}$:
@@ 5.1.3. Pairing Considerations
Selection of the binding modality. We can observe that the quality of the pseudo-paired dataset depends heavily on whether features of the shared modality can accurately measure the semantic similarity of different data samples. Therefore, when the disparate datasets $D_{A}$ and $D_{B}$ have multiple overlapping modalities, it is crucial to carefully select the shared modality $\mathbf{m}_{s}$ for data binding.
Here, we summarize two major properties that a good shared modality should possess for data pairing in MMBind. First, the shared modality itself should be able to effectively distinguish different events in the downstream task. Otherwise, the features generated by the unimodal encoder will fail to differentiate between similar and dissimilar events. For example, in the task of classifying an object’s color, mmWave radar is not an appropriate choice for a common modality, as the similarity measured in the latent space would not reflect the object’s color. Using this mismatched multimodal data for training is likely to degrade the final performance.
@@
Binding more than two datasets. The above approach can be easily extended to integrate multiple incomplete datasets through successive data binding, even when utilizing different shared modalities. However, while incorporating pseudo-paired data typically enhances performance compared to using no pairings, the improvement does not always correlate with the number of paired samples. When integrating multiple datasets, it is essential to consider their heterogeneity. For instance, there may be substantial differences in data distribution among the paired datasets due to reasons such as variations in sensor placement. We also evaluate and discuss the impact of additional paired data in Section 6.3.
@@ 6.1. Methodology
6.1.1. Datasets
We evaluate the performance of MMBind using ten real-world multimodal datasets with various data modalities and domain shifts. Table 2 summarizes the datasets and their data splitting. We use the UTD (Chen et al., 2015), MM-Fi (Yang et al., 2024), PAMAP2 (Reiss and Stricker, 2012), and SUN-RGBD  (Song et al., 2015; Silberman et al., 2012; Janoch et al., 2013; Xiao et al., 2013) (subset) datasets to simulate cross-node (subject) data pairing. We use MotionSense (Malekzadeh et al., 2019), Shoaib (Shoaib et al., 2014) and RealWorld (Sztyler and Stuckenschmidt, 2016) datasets, and GR4DHCI (Wang et al., 2023), DHG (Devineau et al., 2018) and Briareo (Manganaro et al., 2019) dataset to evaluate the performance of MMBind in cross-dataset data binding. When performing cross-node data binding, we construct two unlabeled sub-datasets $\{\mathbf{m}_{1},\mathbf{m}_{s}\}$ and $\{\mathbf{m}_{2},\mathbf{m}_{s}\}$, where $\mathbf{m}_{s}$ can be either a shared sensor modality or the data label. A small modal-complete, labeled dataset $\{\mathbf{m}_{1},\mathbf{m}_{2}\}$ is used for supervised finetuning, while a larger dataset with the same modalities is used for testing. For sensor data as binding modality, we use Acc as the shared modality for UTD-MHAD, PAMAP2, Skeleton for MM-Fi, and Image for SUN-RGBD. Details on cross-dataset binding are in Section 6.3.
@@ 6.1.3. Configurations
For the unimodal feature encoders, we use CNN+RNN layers or transformers to extract deep features. Specifically, 2D-CNN is applied to inertial data and image-like data, 3D-CNN to skeleton data, transformers to radar data, and multi-layer perceptrons are used for the classifier. We employ an encoder-only LLM, paraphrase-MiniLM-L6-v2 trained by Sentence-BERT (Reimers, 2019), to measure the similarity of language labels. Each experiment is repeated five times.
@@ 6.2.3. Label as the binding modality
6.3. Performance on Cross-dataset Binding
In this section, we evaluate the performance of MMBind in cross-dataset binding.
6.3.1. Performance in different applications
As discussed in Section 3, we can bind data from two different datasets to enhance the performance on a third distinct dataset (after finetuning on the third). We evaluate this under two applications. For human activity recognition, we bind data (i.e., (Acc, Gyro, Mag)) from the MotionSense and Shoaib-right datasets, then finetune/test on RealWorld. We also test gesture recognition by binding data (i.e., (Skeleton, Depth, IR)) from GR4DHCI and DHG, then finetune/test on Briareo. Figure 9 and 10 showcase the performance on RealWorld and Briareo, respectively, with different proportions of finetuning data. We observe consistent improvements from adding pseudo-paired data across both label and sensor binding, particularly when finetuning with limited data. For example, on the RealWorld dataset, pre-training the model with pseudo label-paired data samples significantly boosts performance with only 1% training data (77.9%), which approaches the performance of using 10% training data (79.1%).
6.3.2. Impact of pairing more than two datasets.
Table 5 illustrates the performance of binding more than two datasets. While adding pseudo-paired data typically enhances performance compared to using no pairing, the improvement does not always correlate with the number of paired samples. For instance, when incorporating the Shoaib-left dataset, accuracy increases by 2.39% with an additional 4,500 paired samples. However, adding paired data from the Shaib-wrist dataset results in a slight drop in performance. This may be attributed to the significant differences in data distribution between the Shaib-wrist and Shaib-right datasets, which arise from variations in sensor placement. Therefore, when pairing multiple datasets, it is essential to consider the heterogeneity of the different paired datasets.
@@ 6.5.2. Different pairing schemes
During data pairing, we can either pair each sample with the one having the highest feature similarity or with multiple samples of lower similarity. We explore this trade-off between the size of the pseudo-paired dataset (more paired data) and the pairing accuracy (more accurately paired data) using the MM-Fi dataset. The 432 samples were obtained by pairing most similar samples, while the 169 and 2,610 samples were obtained by setting various thresholds on similarity. Table 7 shows that downstream accuracy will suffers at both extremes. Therefore, in MMBind, we choose the scheme of highest-similarity pairing (i.e., 432 samples).
@@ 6.6.1. Training overhead on edge
We now evaluate the robustness of the multimodal embedding trained by MMBind.
Robustness to different environments. We evaluate the performance of MMBind when deployed in different environments during inference. First, using the MM-Fi dataset, we enforce that the two pre-training incomplete datasets and testing dataset all contain unique, non-overlapping environments. Despite these variations, MMBind outperforms all baselines, achieving an 81.65% classification accuracy—over 8% higher than the best baseline. Moreover, as detailed in Section 6.3, MMBind performs well when binding different datasets with various protocols, subjects, and environments.
Robustness to different modality combinations. Table 8 shows the performance of MMBind when deployed on various modality combinations during inference. The model is pre-trained with paired data from (Skeleton, Gyro) and original (Acc, Gyro) and (Acc, Skeleton) data, and fine-tuned on limited labeled data with different modality combinations. In addition to the data combinations used during pre-training, MMBind significantly improves over the lower bound on unseen modality combinations, such as unimodal data (Acc, Gyro, Skeleton) and fully-modality (Acc + Gyro + Skeleton). This indicates that the multimodal embeddings learned by MMBind are robust across different modality combinations.
@@ Acknowledgements.
The research reported in this paper was sponsored in part by the DEVCOM Army Research Laboratory under Cooperative Agreement #W911NF-17-2-0196, the National Institutes of Health under Award #1P41EB028242, and the Air Force Office of Scientific Research under Cooperative Agreement #FA95502210193. Any findings in this material are those of the author(s) and do not reflect the views of any of the above funding agencies. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation here on.

### v2

@@ 5.1. Pairing Incomplete Data with Shared Modality
To enhance the effectiveness and scalability of data pairing, our key idea is to measure the similarity of feature embeddings of the shared modality data rather than comparing high-dimensional raw sensor data. Therefore, we train a unimodal feature encoder via reconstruction to compress the shared modality data into a latent space. Such a latent space is more structured than the raw data distributions, where semantically similar samples are grouped closely together, while dissimilar samples are positioned further apart (Connor et al., 2021; Kirchoff et al., 2024; Xu et al., [n. d.]).
@@
Similarly, when labels serve as the shared modality, we can directly utilize pre-trained large language models (LLMs) to measure the semantic similarity of natural language labels, eliminating the need to train unimodal encoders. For example, we employ an encoder-only LLM, paraphrase-MiniLM-L6-v2, trained by Sentence-BERT (Reimers, 2019), to measure the similarity of language labels. In this paper, we primarily focus on cases where labels sufficiently describe the similarity of the data. However, class labels can sometimes fail to fully capture the similarity of an event, leading to incongruent data associations. For example, sensor data labeled identically but collected from different locations (e.g., worn inwards versus outwards) may exhibit significant differences. In such cases, we can enhance the labels by incorporating additional metadata in structured formats, such as the subjects’ gender, environment, and devices used, alongside the class labels. The LLM then identifies the most semantically similar samples and calculates their corresponding similarities based on the language tokens of the labels.
@@ 5.1.2. Generating Pseudo-Paired Data
As mentioned earlier, when labels serve as the shared modality, we utilize existing pre-trained LLMs to tokenize the language labels and derive their corresponding embeddings $\mathbf{h}^{j}_{s}$ and $\mathbf{h}^{k}_{s}$. Next, we compute the pairwise cosine similarity $a_{jk}$ between features of the shared modality $\mathbf{h}^{j}_{s}$ from $D_{A}$ and $\mathbf{h}^{k}_{s}$ from $D_{B}$, resulting in a similarity matrix $\mathbf{A}\in\mathbb{R}^{|D_{A}|\times|D_{B}|}$. We use pairwise cosine similarity as the pairing metric, as it is a common method for measuring the similarity of feature representations in deep learning models. However, other similarity measurement metrics, such as Euclidean distance, can be also employed in our design.
@@
Here, $\arg\max$ denotes the index where $a_{jk}$ achieves its maximum value within $\mathbf{a}_{j}$. Therefore, the corresponding paired sample from $D_{B}$ will be $\mathbf{X}^{p_{j}}_{B}=(\mathbf{m}^{p_{j}}_{2},\mathbf{m}^{p_{j}}_{s})$. Moreover, each pseudo-paired multimodal data sample will also be associated with the corresponding similarity measurement $a_{jp_{j}}$. This process occurs for every $j_{th}$ data sample in $D_{A}$ and for every $k_{th}$ data sample in $D_{B}$:
@@
When pairing large datasets or when feature embeddings of shared modalities have high dimensionality, pairwise binding by exhausting the cosine similarity between every pair of samples can be highly time-consuming. To address this, we could explore group-based pairing to reduce the number of data samples in search of optimal pairs, or employ techniques such as Principal Component Analysis (PCA) (Abdi and Williams, 2010) to reduce the dimensionality of features, further accelerating the data pairing process.
@@ 5.1.3. Pairing Considerations
Selection of the binding modality. The quality of the pseudo-paired dataset depends heavily on whether features of the shared modality can accurately measure the semantic similarity of different data samples. As a result, the performance of MMBind heavily relies on the shared modalities. Therefore, when the disparate datasets $D_{A}$ and $D_{B}$ have multiple overlapping modalities, it is crucial to carefully select the shared modality $\mathbf{m}_{s}$ for data binding.
Here, we summarize two major properties that a good shared modality should possess for data pairing in MMBind. If the modality does not exhibit such properties, there may be no performance benefits when binding different datasets, as shown in Section 6.5.1. First, the shared modality itself should be able to effectively distinguish different events in the downstream task. Otherwise, the features generated by the unimodal encoder will fail to differentiate between similar and dissimilar events. For example, in the task of classifying an object’s color, mmWave radar is not an appropriate choice for a common modality, as the similarity measured in the latent space would not reflect the object’s color. Using this mismatched multimodal data for training is likely to degrade the final performance.
@@
Binding more than two datasets. The above approach can be easily extended to integrate multiple incomplete datasets through successive data binding. When there are multiple datasets, the shared modality does not need to be the same across different datasets, as evaluated in Section 6.3.3. However, while incorporating pseudo-paired data typically enhances performance compared to using no pairings, the improvement does not always correlate with the number of paired samples. When integrating multiple datasets, it is essential to consider their heterogeneity. For instance, there may be substantial differences in data distribution among the paired datasets due to reasons such as variations in sensor placement. We also evaluate and discuss the impact of additional paired data in Section 6.3.
@@ 6.1. Methodology
6.1.1. Datasets
We evaluate the performance of MMBind using ten real-world multimodal datasets with various data modalities and domain shifts. Table 2 summarizes the datasets and their data splitting. We use the UTD (Chen et al., 2015), MM-Fi (Yang et al., 2024), PAMAP2 (Reiss and Stricker, 2012), and SUN-RGBD  (Song et al., 2015; Silberman et al., 2012; Janoch et al., 2013; Xiao et al., 2013) (subset) datasets to simulate cross-node (subject) data pairing. We use MotionSense (Malekzadeh et al., 2019), Shoaib (Shoaib et al., 2014) and RealWorld (Sztyler and Stuckenschmidt, 2016) datasets, and GR4DHCI (Wang et al., 2023), DHG (Devineau et al., 2018) and Briareo (Manganaro et al., 2019) dataset to evaluate the performance of MMBind in cross-dataset data binding. When performing cross-node data binding, we construct two unlabeled sub-datasets $\{\mathbf{m}_{1},\mathbf{m}_{s}\}$ and $\{\mathbf{m}_{2},\mathbf{m}_{s}\}$, where $\mathbf{m}_{s}$ can be either a shared sensor modality or the data label. A small modal-complete, labeled dataset $\{\mathbf{m}_{1},\mathbf{m}_{2}\}$ is used for supervised finetuning, while a larger dataset with the same modalities is used for testing. For sensor data as binding modality, we use Acc as the shared modality for UTD-MHAD, PAMAP2, Skeleton for MM-Fi, and Image for SUN-RGBD. Details on cross-dataset binding are in Section 6.3. Among these datasets, PAMAP2, SUN-RGBD, RealWorld, MotionSense, GR4DHCI, and Briareo are collected under uncontrolled, natural, and real-world settings, exhibiting significant domain shifts across different subjects and environments.
@@ 6.1.3. Configurations
For the unimodal feature encoders, we use CNN+RNN layers or transformers to extract deep features. Specifically, 2D-CNN is applied to inertial data and image-like data, 3D-CNN to skeleton data, transformers to radar data, and multi-layer perceptrons are used for the classifier. We employ an encoder-only LLM, paraphrase-MiniLM-L6-v2 trained by Sentence-BERT (Reimers, 2019), to measure the similarity of language labels. We repeat each experiment on five different seeds.
@@ 6.2.3. Label as the binding modality
6.3. Dealing with Domain Shift
In this section, we evaluate the performance of MMBind under substantial domain shift.
@@ 6.3. Dealing with Domain Shift
6.3.1. Performance in cross-dataset binding.
As discussed in Section 3, we can bind data from two different datasets to enhance the performance on a third distinct dataset (after finetuning on the third). We evaluate this under two applications. For human activity recognition, we bind data (i.e., (Acc, Gyro, Mag)) from the MotionSense and Shoaib-right datasets, then finetune/test on RealWorld. We also test gesture recognition by binding data (i.e., (Skeleton, Depth, IR)) from GR4DHCI and DHG, then finetune/test on Briareo. Figure 9 and 10 showcase the performance on RealWorld and Briareo, respectively, with different proportions of finetuning data. We observe consistent improvements from adding pseudo-paired data across both label and sensor binding, particularly when finetuning with limited data. For example, on the RealWorld dataset, pre-training the model with pseudo label-paired data samples significantly boosts performance with only 1% training data (77.9%), which approaches the performance of using 10% training data (79.1%). Therefore, MMBind performs well when binding different datasets with various protocols, subjects, and environments.
6.3.2. Performance with different levels of domain shifts.
In Table 5, we showcase MMBind’s performance in the presence of varying domain shifts between the two training datasets that are bound together. We utilize the different environments naturally present in the MM-Fi dataset to emulate various levels of domain shift. In the case of “No Domain Shift”, each binding dataset contains an equal number of unique subjects from Environment 1 and 2. With “Medium Domain Shift”, the first dataset will have 80% of its subjects from Environment 1, with the rest from Environment 2, with a similar allocation for the second dataset. Finally, with “Large Domain Shift”, all subjects in the first dataset are from Environment 1, and all subjects in the second dataset are from Environment 2. We can see that MMBind retains its advantage over the baselines regardless of domain shift, exemplifying the utility of this approach.
Domain Shift Lower Bound Unimodal MIM MPM CMG DCM Image Bind MMBind Upper Bound
None 68.43 33.55 72.94 64.46 73.21 71.72 73.76 79.48 78.68
Medium 67.69 55.58 72.30 72.46 72.66 73.81 72.61 78.48 75.14
Large 68.13 62.41 66.11 76.00 75.81 72.57 72.25 79.40 78.37
Table 5. Accuracy under varying domain shifts.
@@
6.3.3. Impact of pairing more than two datasets.
Table 6 illustrates the performance of binding more than two datasets. While adding pseudo-paired data typically enhances performance, the improvement does not always correlate with the number of paired samples. For instance, binding data from the left pocket and right pocket improves performance (MotionSense + Shoaib-right + Shoaib-left), while binding data from the wrist and right pocket results in a slight performance drop (MotionSense + Shoaib-right + Shoaib-wrist). This may be attributed to the significant differences in data distribution between the Shaib-wrist and Shaib-right datasets, which arise from variations in sensor placement. Therefore, MMBind is effective when sensor locations are similar and do not affect the data distributions of the same events. Moreover, when pairing multiple datasets, it is essential to consider the heterogeneity of the different paired datasets.
@@ 6.5.2. Different pairing schemes
During data pairing, we can either pair each sample with the one having the highest feature similarity or with multiple samples of lower similarity. We explore this trade-off between the size of the pseudo-paired dataset (more paired data) and the pairing accuracy (more accurately paired data) using the MM-Fi dataset. The 432 samples were obtained by pairing most similar samples, while the 169 and 2,610 samples were obtained by setting various thresholds on similarity. Table 8 shows that downstream accuracy will suffers at both extremes. Therefore, in MMBind, we choose the scheme of highest-similarity pairing.
@@ 6.6.1. Training overhead on edge
We now evaluate the robustness of the multimodal embedding trained by MMBind. Table 9 shows the performance of MMBind when deployed on various modality combinations during inference. The model is pre-trained with paired data from (Skeleton, Gyro) and original (Acc, Gyro) and (Acc, Skeleton) data, and fine-tuned on limited labeled data with different modality combinations. In addition to the data combinations used during pre-training, MMBind significantly improves over the lower bound on unseen modality combinations, such as unimodal data (Acc, Gyro, Skeleton) and fully-modality (Acc + Gyro + Skeleton). This indicates that the multimodal embeddings learned by MMBind are robust across different modality combinations.
@@ Acknowledgements.
The research reported in this paper was sponsored in part by the DEVCOM Army Research Laboratory under award #W911NF-17-2-0196, the National Institutes of Health under award #1P41EB028242, and the Air Force Office of Scientific Research under awards # FA95502210193 and FA95502310559. Any findings in this material are those of the author(s) and do not reflect the views of any of the above funding agencies. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation here on.